OpenStudy (ronaldo7):

A chemistry intern needs 20 liters of a 15% sodium solution. He decided to mix a 10% solution with a 30% solution. How many liters of a 10% solution and a 30% solution does he need.

OpenStudy (ronaldo7):

may I ask how did you set that up please

OpenStudy (anonymous):

its wrong

OpenStudy (ronaldo7):

So how is it wrong levinelevine no problem

OpenStudy (anonymous):

im solving it, is it 20% in mass or in mole? or it doesnt say?

OpenStudy (ronaldo7):

the question is all I have for in formation

OpenStudy (anonymous):

ok, ive got it. here you go

OpenStudy (anonymous):

first. we have a 2 equation with 2 variable system. the fist equation can be determined (x+y)=20 where x is the liters of solution 10% and y 30%. then the second equation is 0.1*X+0.3Y=20*0.15. solving this two equation will give the answer. do you know how to do this?

OpenStudy (anonymous):

what was ur answer @AIK ? I got 5liter of 30% and 15 of 10%

OpenStudy (ronaldo7):

hold on please

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well clearly the intern is unsophisticated and does not know what he is doing

OpenStudy (anonymous):

that is what I got, so I was right after all. Just was not sure since I am pretty tired and did it in my head half dead.

OpenStudy (ronaldo7):

what us .15*20

OpenStudy (ronaldo7):

is*

OpenStudy (anonymous):

its the 20 liters of Na multiplied by its concentration in fraction (15%/100)=0.15

OpenStudy (anonymous):

Dude levin you are smart

OpenStudy (anonymous):

thanx @Barcelona20012 :)

OpenStudy (anonymous):

Excuse me Ronaldo Barcelona is clearly better than Real Madrid therefore you may want to change your name

OpenStudy (ronaldo7):

levine i am confused here is what i got so far